# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 8.5. The Hacking–Keel–Yu conjecture
Hence, it remains to treat the case $I=J$, that is, when $\mathcal{Z}_{\sigma}\rightarrow\mathcal{Z}_{\tau}$ is an isomorphism in codimension one. If $\mathcal{Z}_{\sigma}\rightarrow\mathcal{Z}_{\tau}$ contracts a curve in the double locus of $\mathcal{Z}_{\sigma}$, then the corresponding edge in $\mathscr{P}_{H_{\sigma}}$ disappears in $\mathscr{P}_{H_{\tau}}$, and so $Y_{\sigma}$, and $Y_{\tau}$ are different. Finally, if $\mathcal{Z}_{\sigma}\rightarrow\mathcal{Z}_{\tau}$ does not contract any curve in the double locus of $\mathcal{Z}_{\sigma}$, then the exceptional locus of $\mathcal{Z}_{\sigma}\rightarrow\mathcal{Z}_{\tau}$ is a rigid internal curve in an irreducible component of $\mathcal{Z}_{\tau}$, whose strict transform in $\mathcal{X}$ is necessarily an internal exceptional curve. It follows that $NE(\mathcal{Z}_{\tau}/\overline{\mathcal{X}}_{\mathscr{P}})$ is contained in the codimension one face of $NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}}_{\mathscr{P}})$ defined by the M1 flop induced by this internal exceptional curve, which is by Lemma 6.19 in contradiction with the assumption that $\tau$ is a codimension cone of the secondary fan. Hence, this last case actually does not occur, and this concludes the proof that the map $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\to\mathcal{M}_{(Y,D,L)}$ is finite.
@@
To complete the proof of Theorem 8.11, it remains to show that the map $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\to\mathcal{M}_{(Y,D,L)}$ is surjective. The moduli stack of log Calabi–Yau surfaces deformation equivalent to $(Y,D)$ is irreducible of dimension $Q-2$, where $Q$ is the charge of $(Y,D)$ as in [34, Definition 1.1]. Indeed, fixing a toric model $(\overline{Y},\overline{D})$, this space is covered by the torus $(\mathbb{C}^{\star})^{Q}$ of positions of the points to blow-ups on $\overline{D}$, quotiented by the action of $(\mathbb{C}^{\star})^{2}$ by toric automorphisms. Moreover, given such a log Calabi–Yau pair $(Y^{\prime},D^{\prime})$, the space of curves $C^{\prime}\in|L^{\prime}|$ such that $(Y^{\prime},D^{\prime}+\epsilon\,C^{\prime})$ is a KSBA stable pair, is open, and so dense in $|L^{\prime}|$ if non-empty. Hence, the space of KSBA stable pairs deformation equivalent to $(Y,D,L)$ is irreducible, and so the closure $\mathcal{M}_{(Y,D,L)}$ of this space in the moduli space of all KSBA stable pairs is also irreducible. On the other hand, the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ is complete, and so the toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is proper. Therefore, it is enough to prove that the map $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\to\mathcal{M}_{(Y,D,L)}$ is dominant.
@@
Fix a large enough positive integer $k$ and a refinement $\widetilde{\mathscr{P}}$ of the scaled polyhedral decomposition $kP$ as in Lemma 8.3. By Theorem 8.10, the general fiber $(Y_{t},D_{t})$ of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is deformation equivalent to $(Y,D)$, and so admits a toric model $(Y_{t},D_{t})\leftarrow(\widetilde{Y}_{t},\widetilde{D}_{t})\rightarrow(\overline{Y},\overline{D})$, where $(\overline{Y},\overline{D})$ is the toric variety with momentum polytope $\overline{P}$, and the exceptional curves $E_{ij}$ of $\widetilde{Y}_{t}\rightarrow\overline{Y}$ are in one-to-one correspondence with the cut triangles $\Delta_{ij}$. Moreover, it follows from the proof of Theorem 8.10 that $(\widetilde{Y}_{t},\widetilde{D}_{t})$ admits an almost toric fibration over $(kP)^{gs}$, and that the classes of the exceptional curves $E_{ij}$ are the classes in $\widetilde{Y}_{t}$ obtained by lifting as in [70, §6.2], [82] the tropical 1-cycles defined by the cuts $\delta_{ij}$ ending at the focus-focus singularities $\widetilde{p}_{ij}$ as in §8.2. In particular, the period calculation proving [70, Proposition 6.15], based on the main result of [82], directly generalizes to our set-up to show that the marked period of $(Y_{t},D_{t})$ marked by $\mathcal{O}_{Y_{t}}(E_{ij})$, as defined in [41], is equal to $t^{E_{ij}}$. This shows that every log Calabi–Yau surface deformation equivalent to $(Y,D,L)$ appears as a fiber $(Y_{t},D_{t},L_{t})$ of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$. Finally, given such a fiber $(Y_{t},D_{t},L_{t})$, it follows from the torus action given by Lemma 6.27 that for every $\mathbf{a}:=(a_{p})_{p\in P_{\mathbb{Z}}}\in(\mathbb{C}^{\star})^{P_{\mathbb{Z}}}$, denoting by $C_{\mathbf{a}}$ the curve in $Y_{t}$ defined by the equation $\sum_{p\in P_{\mathbb{Z}}}a_{p}\vartheta_{p}=0$, the pair $(Y_{t},D_{t}+\epsilon\,C_{\mathbf{a}})$ appears as a fiber of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\mathscr{P}}})\rightarrow S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$. Since $(\vartheta_{p})_{p\in P_{\mathbb{Z}}}$ is a basis of $H^{0}(Y_{t},L_{t})$, the set of curves of the form $C_{\mathbf{a}}$ is open dense in the linear system $|L_{t}|$. Hence, one obtains that an open dense subset of $\mathcal{M}_{(Y,D,L)}$ is contained in the image of $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\rightarrow\mathcal{M}_{(Y,D,L)}$, and this concludes the proof of Theorem 8.11. ∎
@@
By Theorem 8.11, $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is a finite cover of the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$, which is projective by [36, 68]. ∎

### 30 Mar 2024 v2

@@ 8.5. The Hacking–Keel–Yu conjecture
Assume first that $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is an isomorphism in codimension one. If $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ contracts a curve in the double locus of $\mathcal{Z}_{1}$, then the corresponding edge in $\mathscr{P}_{H_{1}}$ disappears in $\mathscr{P}_{H_{0}}$. Hence, $Y_{\rho_{1}}$, and $Y_{\rho_{0}}$ are not isomorphic, and the result is proved in this case. On the other hand, if $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ does not contract any curve in the double locus of $\mathcal{Z}_{1}$, then the exceptional locus of $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is a rigid internal curve in an irreducible component of $\mathcal{Z}_{1}$, whose strict transform in $\mathcal{X}^{\prime}$ is necessarily an internal exceptional curve. It follows that $NE(\mathcal{Z}_{0}/\overline{\mathcal{X}}_{\mathscr{P}})$ is contained in the codimension one face of $NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}}_{\mathscr{P}})$ defined by the M1 flop defined by contracting this internal exceptional curve, which is by Lemma 6.19 in contradiction with the assumption that $\rho_{0}$ is a codimension cone of the secondary fan, and so this last case actually does not occur.
Thus, it remains to treat the case where $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is a divisorial contraction. Since $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is of relative Picard rank one, its exceptional locus consists of a single divisor $Z_{1}^{v}$ corresponding to a vertex $v$ of $\mathscr{P}_{H_{1}}$. If an irreducible component of the boundary of $Z_{1}^{v}$ is contracted into $\mathcal{Z}_{0}$, then the corresponding edge of $\mathscr{P}_{H_{1}}$ disappears in $\mathscr{P}_{H_{0}}$, and so we also deduce in this case that $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are not isomorphic. If no boundary component of $Z_{1}^{v}$ is contracted in $\mathcal{Z}_{0}$, then, as in [50, §8.2], $Z_{1}^{v}$ is necessarily a generically $\mathbb{P}^{1}$-bundle over $\mathbb{P}^{1}$, contracted to its base $\mathbb{P}^{1}$ in $\mathcal{Z}_{0}$, and its boundary $\partial Z_{v}^{1}$ is either an irreducible nodal curve mapping 2:1 to the base, or the union of two sections intersecting in two points. If $\partial Z_{v}^{1}$ is the union of two irreducible components $D_{1}$ and $D_{2}$, consider the corresponding edges $E_{1}$ and $E_{2}$ adjacent to $v$ in $\mathscr{P}_{H_{1}}$, and choose a 2-dimensional face $\sigma_{1}$ of $\mathscr{P}_{H_{1}}$ such that $E_{1}\cup E_{2}$ are edges of the boundary of $\sigma_{1}$. Denote by $\sigma_{0}$ the same face viewed as a 2-dimensional face of $\mathscr{P}_{H_{0}}$. By Theorem 7.12, there is an irreducible component $(Y_{\sigma_{1}},\partial Y_{\sigma_{1}})$ (resp. $(Y_{\sigma_{0}},\partial Y_{\sigma_{0}})$) of $Y_{\rho_{1}}$ (resp. $Y_{\rho_{0}}$) isomorphic to the general fiber of the mirror family to the open Kulikov degeneration $\mathcal{X}_{\sigma_{1}}$ (resp. $\mathcal{X}_{\sigma_{0}}$) with tropicalization $(C(\sigma_{1}),C(\mathscr{P}_{\sigma_{1}}))$ (resp. $(C(\sigma_{1}),C(\mathscr{P}_{\sigma_{1}}))$). The vertex $v$ can be viewed as a boundary vertex of both $\mathscr{P}_{\sigma_{1}}$ and $\mathscr{P}_{\sigma_{0}}$. Since the $\mathbb{P}^{1}$-fibers of $Z^{1}_{v}$ are not contracted in $\mathcal{Z}_{1}$, the non-compact irreducible component corresponding to $v$ in the central fiber $\mathcal{X}_{\sigma_{1},0}$ of $\mathcal{X}_{\sigma_{1}}$ is a 00-surface. Hence, by Lemma 6.24, it corresponds to a singular point of $\partial Y_{\sigma_{1}}$. On the other hand, since the $\mathbb{P}^{1}$-fibers of $Z^{1}_{v}$ are contracted in $\mathcal{Z}_{0}$, the non-compact irreducible component corresponding to $v$ in the central fiber $\mathcal{X}_{\sigma_{0},0}$ of $\mathcal{X}_{\sigma_{0}}$ is a $1$-surface. Thus, by Lemma 6.24, it corresponds to a smooth point of $\partial Y_{\sigma_{0}}$. Since a singular point of $\partial Y_{\sigma_{1}}$ becomes a smooth point in $\partial Y_{\sigma_{0}}$, it follows that $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are not isomorphic.
Finally, the case where $\partial Z^{1}_{v}$ is irreducible follows from the previous case as in [50, §8.2]. Indeed, denoting by $\sigma_{1}$ (resp. $\sigma_{0}$) the 2-dimensional face of $\mathscr{P}_{H_{1}}$ (resp. $\mathscr{P}_{H_{0}}$) containing $v$, and cutting $\sigma_{1}$ (resp. $\sigma_{0}$) along the edges of $\mathscr{P}_{\sigma_{1}}$ (resp. $\mathscr{P}_{\sigma_{0}}$) connecting $v$ to the boundary of $\sigma_{1}$ (resp. $\sigma_{0}$), we obtain an integral affine manifold with singularities $(\tilde{\sigma}_{1},\mathscr{P}_{\tilde{\sigma}_{1}})$ (resp. $(\tilde{\sigma}_{0},\mathscr{P}_{\tilde{\sigma}_{0}})$) with the same formal properties as $(\sigma_{1},\mathscr{P}_{\sigma_{1}})$ (resp. $(\sigma_{0},\mathscr{P}_{\sigma_{0}})$) in the case where $\partial Z^{1}_{v}$ is reducible. There is a natural open Kulikov degeneration $\mathcal{X}_{\widetilde{\sigma}_{1}}$ (resp. $\mathcal{X}_{\widetilde{\sigma}_{0}}$) with tropicalization $(C(\widetilde{\sigma}_{1}),C(\mathscr{P}_{\widetilde{\sigma}_{1}}))$ (resp. $(C(\widetilde{\sigma}_{0}),C(\mathscr{P}_{\widetilde{\sigma}_{0}}))$), and the irreducible component $Y_{\tilde{\sigma}_{1}}$ (resp. $Y_{\tilde{\sigma}_{0}}$) of the general fiber of the corresponding mirror family is the normalization of $Y_{\sigma_{1}}$ (resp. $Y_{\sigma_{0}}$). As above, a singular point of $\partial Y_{\tilde{\sigma}_{1}}$ becomes smooth in $\partial Y_{\tilde{\sigma}_{0}}$. It follows that $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are also non-isomorphic in this case, and this concludes the proof that the map $\mathcal{S}_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}\to\mathcal{M}_{(Y,D,L)}$ is finite.
@@
To complete the proof of Theorem 8.11, it remains to show that the map $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to\mathcal{M}_{(Y,D,L)}$ is surjective. The moduli stack of log Calabi–Yau surfaces deformation equivalent to $(Y,D)$ is irreducible of dimension $Q-2$, where $Q$ is the charge of $(Y,D)$ as in [34, Definition 1.1]. Indeed, fixing a toric model $(\overline{Y},\overline{D})$, this space is covered by the torus $(\mathbb{C}^{\star})^{Q}$ of positions of the points to blow-ups on $\overline{D}$, quotiented by the action of $(\mathbb{C}^{\star})^{2}$ by toric automorphisms. Moreover, given such a log Calabi–Yau pair $(Y^{\prime},D^{\prime})$, the space of curves $C^{\prime}\in|L^{\prime}|$ such that $(Y^{\prime},D^{\prime}+\epsilon\,C^{\prime})$ is a KSBA stable pair, is open, and so dense in $|L^{\prime}|$ if non-empty. Hence, the space of KSBA stable pairs deformation equivalent to $(Y,D,L)$ is irreducible, and so the closure $\mathcal{M}_{(Y,D,L)}$ of this space in the moduli space of all KSBA stable pairs is also irreducible. On the other hand, the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ is complete, and so the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ is proper. Therefore, it is enough to prove that the map $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\to\mathcal{M}_{(Y,D,L)}$ is dominant.
@@
Fix a large enough positive integer $k$ and a refinement $\widetilde{\mathscr{P}}$ of the scaled polyhedral decomposition $kP$ as in Lemma 8.3. By Theorem 8.10, the general fiber $(Y_{t},D_{t})$ of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ is deformation equivalent to $(Y,D)$, and so admits a toric model $(Y_{t},D_{t})\leftarrow(\widetilde{Y}_{t},\widetilde{D}_{t})\rightarrow(\overline{Y},\overline{D})$, where $(\overline{Y},\overline{D})$ is the toric variety with momentum polytope $\overline{P}$, and the exceptional curves $E_{ij}$ of $\widetilde{Y}_{t}\rightarrow\overline{Y}$ are in one-to-one correspondence with the cut triangles $\Delta_{ij}$. Moreover, it follows from the proof of Theorem 8.10 that $(\widetilde{Y}_{t},\widetilde{D}_{t})$ admits an almost toric fibration over $(kP)^{gs}$, and that the classes of the exceptional curves $E_{ij}$ are the classes in $\widetilde{Y}_{t}$ obtained by lifting as in [71, §6.2], [83] the tropical 1-cycles defined by the cuts $\delta_{ij}$ ending at the focus-focus singularities $\widetilde{p}_{ij}$ as in §8.2. In particular, the period calculation proving [71, Proposition 6.15], based on the main result of [83], directly generalizes to our set-up and shows that the marked period of $(Y_{t},D_{t})$ marked by $\mathcal{O}_{Y_{t}}(E_{ij})$, as defined in [42], is equal to $t^{E_{ij}}$. This shows that every log Calabi–Yau surface deformation equivalent to $(Y,D,L)$ appears as a fiber $(Y_{t},D_{t},L_{t})$ of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. Finally, given such a fiber $(Y_{t},D_{t},L_{t})$, it follows from the torus action given by Lemma 6.27 that for every $\mathbf{a}:=(a_{p})_{p\in P_{\mathbb{Z}}}\in(\mathbb{C}^{\star})^{P_{\mathbb{Z}}}$, denoting by $C_{\mathbf{a}}$ the curve in $Y_{t}$ defined by the equation $\sum_{p\in P_{\mathbb{Z}}}a_{p}\vartheta_{p}=0$, the pair $(Y_{t},D_{t}+\epsilon\,C_{\mathbf{a}})$ appears as a fiber of $(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\mathscr{P}}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$. Since $(\vartheta_{p})_{p\in P_{\mathbb{Z}}}$ is a basis of $H^{0}(Y_{t},L_{t})$, the set of curves of the form $C_{\mathbf{a}}$ is open dense in the linear system $|L_{t}|$. Hence, we obtain that an open dense subset of $\mathcal{M}_{(Y,D,L)}$ is contained in the image of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}\rightarrow\mathcal{M}_{(Y,D,L)}$, and this concludes the proof of Theorem 8.11. ∎
@@
By Theorem 8.11, $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is a finite cover of the coarse KSBA moduli space $M_{(Y,D,L)}$, which is projective by [36, 69]. ∎